OpenStudy (anonymous):

Write an equation in slope-intercept form for a line that is perpendicular to the given line and that passes through the given point. 40) (4,0); y=-2x 41) (0,2); x=-7 42) (0,2); y= -2x-8 43) (12,8); y=-3x-9 44) (-1,7); y=-1/4x+6 45) (4,-9); -x+2y=10 46) (2,1); -2x+3y=-12 47) (-2,3); y=-5x+1

OpenStudy (anonymous):

Evaluate D(e) = 3e - 1, where e = 2. D(2) = 3(2) - 1 D(2) = 6 - 1 D(2) = 5

OpenStudy (anonymous):

i don't get it...

OpenStudy (anonymous):

@RampagingUbay You still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So an equation that would meet you specifications would have a slope opposite to the given equations and when the x value for the point which it passes through is inputted, the y value of said point is derived. Did that make sense?

OpenStudy (anonymous):

sorry but no...

OpenStudy (anonymous):

i don't know that much english

OpenStudy (anonymous):

Ok, the let's do an example problem. the equation of f(x) = -2x looks like this: the graph of f(x) = 1/2(x) is perpendicular to f(x) = -2x like this:  Since we know that the equation f(x)=1/2x is perpendicular, we just need to add a number until f(4)=0 for f(x)=1/2x. Does this make sense?

OpenStudy (anonymous):

kind of...

OpenStudy (anonymous):

If you want, I could try and google translate my comments into a language you know better

OpenStudy (anonymous):

no its ok

OpenStudy (anonymous):

could you do a problem from the top

OpenStudy (anonymous):

That was problem 40 that I just started.

OpenStudy (anonymous):

Do you know what an inverse is?

OpenStudy (anonymous):

is it the opposite ?

OpenStudy (anonymous):

Sort of. It's the number flipped upside down. So the inverse of 3/1 would be 1/3.

OpenStudy (anonymous):

can you do number 42-45 please, with out words... ?

OpenStudy (anonymous):

I'd rather you be able to do them on your own first. You need to understand this concept because all required advanced math is based on things like this.

OpenStudy (anonymous):

if you do it, i will look at your work and try to do the other ones please... i been stuck on those problems for a very long time

OpenStudy (anonymous):

I'll do 42. y=-2x-8 perpedicular equation = y=1/2x when x=0, y must equal 2: y=1/2(0) = 0 thus, y=1/2(0) + 2 = 2 The equation is y=1/2x+2

OpenStudy (anonymous):

so 43 is y=1/3x-9 y=1/3x+40

OpenStudy (anonymous):

No, 43 would be y=1/3x to y=1/3x+4. where x=12, y=8 so: y=1/3(12)+4 y=4+4 y=8

OpenStudy (anonymous):

where did you put the 8 ?

OpenStudy (anonymous):

1/3(12) is already 4, so all you need to do to get 8 is add another 4

OpenStudy (anonymous):

Binary, so #44 would be y=-4x+6 ?

OpenStudy (anonymous):

Not quite, it is 4x, not -44x for the slope, meaning you need to start from the slope again

OpenStudy (anonymous):

*-4x

OpenStudy (anonymous):

whats with #45. it looks really wierd

OpenStudy (anonymous):

you just need to simplify it to a y=mx+b form. It's just variable solving for y.

OpenStudy (anonymous):

thanks for everything gtg cya

OpenStudy (anonymous):

no problem